A colony of bacteria has an initial population of 1500 bacteria. After 1 hour, the colony has grown to a size of 1750 bacteria. Which of the following is an
exponential function that approximately describes the number of bacteria as a
function of time t hours after the initial observation of 1500 bacteria?

Found 2 solutions by Boreal, solver91311:
Answer by Boreal(15235)

P=Poe{1+r}^n
1750=1500{1+r}^1, since n=1 hour
1.1166=1+r ; divided 1750/1500 and n=1
r=0.1166
An exponential function that is increasing is Starting value*(1+r)^n; r is the rate per n unit time, and n is the number of times. Here, with only one hour, the rate is 0.1166. This does describe the behavior after 1 hour, but after several hours, there may well be better functions that describe the behavior.

Answer by solver91311(24713)



If you are going to ask a question that includes the words "which of the following..." in the stem of the question, you have to include that which follows. Despite what you and many other students seem to believe, none of us are clairvoyant, nor do we have a crystal ball, tarot cards, or chicken bones that we consult to discern the information that you were apparently too lazy to include. Repost.

John

My calculator said it, I believe it, that settles it